OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f over g(−2). −3 3 −1 1

OpenStudy (anonymous):

It means to put the equation F over G when x for G equals -2

OpenStudy (anonymous):

\[\frac{ -3x-3 }{ 2(-2)+1}\]

OpenStudy (anonymous):

so -3times-2

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

where did you get -3*-2?

OpenStudy (anonymous):

beacues u said x for G equals -2

OpenStudy (anonymous):

Right so it would be 2(-2)

OpenStudy (psymon):

Well, you start with g(x) = Notice how there's an x in the parenthesis and thus an x in your function, 2x+1. Well, if something else goes in the parenthesis, like -2 g(-2), this means instead of having x's, you need to have -2 where all the x's are. That's why Bama replaced x with -2 in his example.

OpenStudy (anonymous):

ok

OpenStudy (psymon):

Mhm, so that's why he wrote theproblem the way he did above. I'm sure he can finish explaining :P

OpenStudy (anonymous):

Do you understand where we got the -2 and how we know where to put it?

OpenStudy (anonymous):

Thanks Pysmon:)

OpenStudy (psymon):

Of course ^_^

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Alright so start by simplifying the equation now...

OpenStudy (anonymous):

i got-3

OpenStudy (anonymous):

Right! So now the equation looks like\[\frac{ -3x-3 }{ -3}\]

OpenStudy (anonymous):

so the final answer is 1

OpenStudy (anonymous):

How did you get 1?

OpenStudy (anonymous):

-3x3-3=-9/3 ohhh its -3 as the final answer

OpenStudy (anonymous):

The final answer when you plug in the original equation should be 0. Your close but thats not it...

OpenStudy (anonymous):

how the final answer are -3,3,1,-1,

OpenStudy (anonymous):

it has to be one of those

OpenStudy (anonymous):

Right and it is, but it isnt 1.

OpenStudy (anonymous):

\[\frac{ -3x-3 }{ -3 }=0\]

OpenStudy (anonymous):

so -1

OpenStudy (anonymous):

YUP:)

OpenStudy (psymon):

Dun dun dun! :D

OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)]. 9 3 −9 −3

OpenStudy (psymon):

So as I mentioned before, when something like f(x) changes to f(2), this means that instead of x's in our function, those x's get replaced with 2. Now this is the same thing, but our x's are replaced with g(-6). So it's best if we get an answer for g(-6) first. So g(x) = 2x+1 and instead of x inside of the parenthesis we have -6. This tells us we want to replace x with -6. So doing that gives us 2(-6) + 1. So what would that be?

OpenStudy (anonymous):

-9

OpenStudy (psymon):

That should give us -11 if we do that part. 2(-6) = -12 then + 1 givws us -11. See how?

OpenStudy (anonymous):

GOOD JOB PYSMON:)

OpenStudy (anonymous):

so it is -9

OpenStudy (psymon):

My brain must not be working, I don't see the correct answer listed.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2(-6)-2}=\frac{ 1 }{ -14}\]

OpenStudy (psymon):

Right. I interpreted the problem that way and: \[\frac{ 1 }{ 2(-6) }-2 = \frac{ -25 }{ 12 }\]

OpenStudy (psymon):

I still must be doing something incredibly dumb to not be getting the answer to be any of his choices, lol.

OpenStudy (anonymous):

This is too confusing. Your friend might have wrote it the wrong way...

OpenStudy (psymon):

I think so, because the answer does not match up with his choices, even if you interpret it in both ways.

OpenStudy (anonymous):

Yea...